OpenStudy (anonymous):

Prove: cos(2x)^2 = (1-cos(4x))/2 Help please someone? Or proving cos(2x) = (1-(2x))/2 would be fine too. Help, please.

OpenStudy (anonymous):

Oops, typo. cos(2x) = (1-cos(2x))/2

OpenStudy (anonymous):

Your sure your supposed to proove this?

OpenStudy (anonymous):

Cos (2x) = 1 - 2sin^2x sin^2 x = (1 - cos2x)/2 Now can you prove it?

OpenStudy (anonymous):

That proves the second one, but I still can't prove the first one :/. And yes I am supposed to prove it.

OpenStudy (anonymous):

I am having a difficult time trying to prove the first one. hang on...

OpenStudy (anonymous):

Okay. I tried a lot of different things too, but it just wouldn't add up.

OpenStudy (anonymous):

Sorry...I got close but it didn't match what you have.

OpenStudy (anonymous):

Awe, okay. That's alright, what did you get?

OpenStudy (anonymous):

1 + cos 4x

OpenStudy (anonymous):

divided by 2

OpenStudy (anonymous):

Hmm, I'm wondering. Is \[ 2\cos(2x) = 1-\cos(2x)\] and how did you prove the second one? Because the way I see it that equation can only be true for certain values of x. And because of that not true. You will see it easy if you draw it up.

OpenStudy (anonymous):

The actually problem was wrong. Sorry. The actually problem to prove is: cos(2x)^2 = (1+cos(4x))/2

OpenStudy (anonymous):

The second one was also wrong, it is supposed to be cos(2x) = (1+(2x))/2

OpenStudy (anonymous):

But you will have the same problem, with the second one. cos(2x) is a periodic function and (1+2x)/2 is a linear function. And this is not always going to be equal either. If you put the equations in in wolfram you will see that none of them hold up.